# Constraining color-charge effects of partonic energy loss with jet axis-based inclusive jet substructure measurement

### v1

@@ 1 Introduction
Ultra-relativistic heavy-ion collision experiments aim to recreate a deconfined state of quarks and gluons, known as the quark-gluon plasma (QGP), under laboratory conditions to investigate the properties of this unique form of nuclear matter. Jets, collimated streams of particles produced in high-energy collisions, serve as versatile tools for probing the QGP Apolinario:2022vzg ; Cunqueiro:2021wls . These jets are initiated by energetic quarks and gluons (partons) produced in high-momentum-transfer processes during the initial stages of the collision. The partons subsequently fragment into lower-energy partons, which then hadronize into final-state particles. As the outgoing partons traverse the QGP, they lose energy through interactions with the medium. These interactions result in medium-induced modifications to various experimental observables, collectively referred to as “jet quenching”. Experimental evidence for jet quenching has been observed using multiple probes at both the BNL RHIC PHENIX:2001hpc ; STAR:2002ggv and the CERN LHC CMS:2021vui ; CMS:2021nhn ; ATLAS:2023hso ; ALICE:2023waz . Such probes include the suppression of high transverse momentum ($p_{\text{T}}$) particle yields, modifications to jet fragmentation functions, and changes in jet shapes. Some recent reviews on this subject can be found in Refs. CMS:2024krd ; ALICE:2022wpn ; Apolinario:2022vzg ; Cunqueiro:2021wls .
@@
In parallel with experimental studies of quenching phenomena, significant theoretical advances have been made over the past two decades in characterizing partonic energy loss in the QGP medium. The mechanisms understood as the primary contributors to the energy loss are medium-induced gluon radiation and collisional energy loss. Despite substantial progress on both theoretical and experimental fronts, the details of parton-medium interactions and the precise mechanisms governing parton energy loss remain unclear. Emerging techniques, such as jet substructure studies and the tagging (or selection) of quark and gluon jets, offer new insights into energy loss mechanisms Kogler:2018hem ; Vertesi:2024tdv . Due to differences in Casimir color factors ($\text{C}_{\text{F}}=4/3$ for quarks and $\text{C}_{\text{A}}=3$ for gluons), gluon-initiated jets are expected to interact more strongly with the QGP and lose more energy compared to quark-initiated jets. This enhanced energy loss reduces the fraction of gluon jets within the surviving jet population in heavy-ion collisions at a given jet $p_{\text{T}}$ compared to the vacuum case without a medium. Recent measurements suggest that the relative fractions of quark and gluon jets are modified in heavy-ion collisions compared to elementary proton-proton collisions, reflecting the differential energy loss due to the distinct color charges of quarks and gluons mediumqg ; eg:partonloss ; ALICE:2019ykw .
@@
The CMS experiment attempted to study gluon-like jet fractions by applying the template-fit method to the jet-charge observable for an inclusive jet sample from lead-lead (PbPb) collisions at a center-of-mass energy per nucleon pair ($\sqrt{s_{{}_{\text{NN}}}}$) of 5.02 TeV CMS:2020plq . This measurement found no significant evidence of a decrease in gluon jets or a corresponding increase in quark jets. The jet-charge observable currently lacks theoretical predictions for the expected in-medium modifications. The ATLAS experiment measured the jet nuclear modification factor ($R_{\text{AA}}$) and estimated the fractional energy loss ($S_{\text{loss}}$) for photon-tagged jets, and compared these observables with those for inclusive jets ATLAS:2023iad . Photon-tagged jets are more likely to be quark-initiated, whereas inclusive jets at low transverse momentum ($p_{\text{T}}<200$ GeV) are predominantly gluon-initiated at the LHC ATLAS:2023iad . It was reported that the $R_{\text{AA}}$ of photon-tagged jets is significantly higher, and their estimated $S_{\text{loss}}$ significantly lower, than those of inclusive jets, providing evidence that quark jets lose less energy than gluon jets in the medium. However, a detailed description of color-charge dependent energy loss in the QGP medium, along with the relative modifications of quark and gluon jet fractions, remains elusive and warrants further experimental and theoretical investigation through additional observables.
@@
Recently, the CMS experiment measured a new jet substructure observable, the jet axis decorrelation ($\Delta\text{j}$), for a similar data sample (inclusive jets from 5.02 TeV PbPb collisions) CMS:2024koh . This observable is predicted to be sensitive to the parton-medium interactions and provide constraints for energy loss models angelJetaxis . The study found that $\Delta\text{j}$ is narrower than predictions from various unquenched models, such as pythia pythia82 and herwig Sirunyan:2020pqv . This narrowing could be attributed to modifications in the internal structure of jets caused by interactions with the QGP medium, selection biases favoring less-quenched jets, and/or results from the color-charge dependence of partonic energy loss. In the latter scenario, the larger average energy loss shifts gluon jets further down in energy compared to quark jets, resulting in a different partonic composition of quenched and unquenched jet samples.
@@
In this paper, we attempt to interpret the recent CMS jet axis decorrelation measurements by comparing them to pythiapythia82 simulations with varying relative quark/gluon fractions and emulated color-charge dependent jet energy loss. A template-fit approach is employed to estimate the lower limit on gluon jet fractions within the inclusive jet sample. The quark- and gluon-initiated jet templates are derived from inclusive jet samples in 5.02 TeV pp collisions simulated with the pythia8 event generator. Simultaneously, the average jet energy loss for inclusive, quark, and gluon jets is estimated by minimizing the chi-square $(\chi^{2})$ between the $\Delta\text{j}$ distributions measured by CMS and those generated by pythia8 with an applied $p_{\text{T}}$ shift imitating energy loss (in this simplistic approximation, we assume all jets in a given sample lose the same fractional amount of energy, e.g., each jet $p_{\text{T}}$ is shifted down by the same percentage).
@@
The paper is organized as follows. Section 2 outlines the analysis framework used to measure jet axis decorrelation. Section 2.1 discusses the template fits, while Section 2.2 focuses on estimating the $p_{\text{T}}$ shifts for inclusive, quark, and gluon jets to interpret the CMS $\Delta\text{j}$ distributions. Section 3 presents the results, followed by a detailed discussion. Finally, Section 4 summarizes the main conclusions of this work.
2 Analysis method and framework
@@
The jet axis decorrelation, $\Delta\text{j}$, is the angular difference between the axes determined by the anti-$k_{\text{T}}$ energy-weighted (E-scheme) recombination and the winner-take-all (WTA) recombination schemes angelJetaxis ; CMS:2024koh . It is defined as
$\Delta\text{j}=\sqrt{(\eta^{\text{E-scheme}}-\eta^{\text{WTA}})^{2}+(\phi^{\text{E-scheme}}-\phi^{\text{WTA}})^{2}},$ (1)
@@
Where $\eta^{\text{E-scheme}}$ ($\eta^{\text{WTA}}$) and $\phi^{\text{E-scheme}}$ ($\phi^{\text{WTA}}$) denote the pseudorapidity and azimuthal angle of the E-scheme (WTA) axis, respectively, which characterize the jet’s direction. In the E-scheme jet recombination Cacciari:2011ma , the jet axis is determined by the four-vector sum of its constituent particles, representing a momentum-averaged direction of energy flow that includes contributions from soft particles. This method clusters jets by iteratively combining particle pairs into pseudo-jets, with the direction of each new pseudo-jet defined by the vector sum of the four-momenta of its constituents. The final four-vector defines the E-scheme jet axis in $(\eta,\phi)$ coordinates. Conversely, the WTA scheme emphasizes the hardest prong at each clustering step, making it more sensitive to collinear radiation. To determine the WTA axis, the jet is reclustered using the WTA algorithm wta1 ; wta2 with the same constituents, where the direction of each new pseudo-jet aligns with the highest-$p_{\text{T}}$ particle or pseudo-jet. This typically results in the WTA axis aligning with the direction of the hardest constituent. The measured $\Delta\text{j}$ distributions are typically normalized to the total jet number in each $p_{\text{T}}$ interval, constituting a shape-only measurement.
@@
The pythia v8.226 pythia82 Monte Carlo event generator, configured with the CP5 tune pythiaCP5tune and NNPDF3.1 parton distribution functions at next-to-next-to-leading order nnpdf31 , is used to simulate particle production in pp collisions. Following the CMS data analysis, final-state jets are reconstructed using the anti-$k_{\text{T}}$ algorithm Cacciari:2008gp with a distance parameter of $R=0.4$, as implemented in the FastJet framework Cacciari:2011ma . To match the acceptance of the experimental work, simulated jets are selected within the kinematic range of $300120<p_{\text{T}}<$ 300 GeV and $|\eta|<1.6$.
@@
In pythia, jets can be identified (or tagged) based on the type of parton originating the shower, allowing differentiation between quark and gluon jets. According to pythia8 simulations, the inclusive jet sample in the transverse momentum range of this study contains nearly equal amounts of quark and gluon jets without any medium effects. The predicted color-charge dependence of partonic energy loss in the presence of QGP is expected to result in a more significant migration of gluon-initiated jets towards lower final-state energy values than quark jets. Since gluon jet $\Delta\text{j}$ distributions are significantly broader than those of quark jets, differences in energy loss (and corresponding migration rates) could also contribute to the narrowing of the jet-axis decorrelations observed in PbPb data by the CMS experiment CMS:2024koh .
2.1 Template fits
@@
A template fit was performed on the CMS PbPb data using distributions associated with different partons derived from pythia simulations. The primary goal was to investigate a potential shift in the rates between quark and gluon jets. Since pythia is a vacuum-like (unquenched) simulation, this fitting attempt is designed to capture the full extent of the observed modifications in the CMS PbPb data by reducing the relative abundance of gluon jets in the measured inclusive jet sample. All quark flavors are combined into a single quark template, and a two-component quark-gluon template fit is applied to each measured CMS distributions. Figures 1-4 illustrate the template fits for different centrality and jet $p_{\text{T}}$ intervals. Centrality is defined as the fraction of the total nucleus-nucleus cross section, with 0% corresponding to the maximum overlap of the colliding nuclei. The fit range is shown in the $\Delta\text{j}$ distributions, and the lower panels display the ratio of data to fits.
@@
Figure 1: Template fits for the inclusive jet $\Delta\text{j}$ distributions in PbPb collisions at 50–80% centrality (black circles), measured by the CMS experiment, across different $p_{\text{T}}$ intervals. The quark and gluon jet templates from pythia are represented by the red and blue shaded areas, respectively. The lower panels show the ratio of data to the fits. For the CMS data points, systematic uncertainties are indicated by rectangles (correlated uncertainties) and shaded areas (uncorrelated uncertainties), while vertical solid lines depict statistical uncertainties. In the lower panels, the vertical lines represent the total uncertainties in the data.
Figure 2: Template fits for the inclusive jet $\Delta\text{j}$ distributions in PbPb collisions at 30–50% centrality (black circles), measured by the CMS experiment, across different $p_{\text{T}}$ intervals. The quark and gluon jet templates from pythia are represented by the red and blue shaded areas, respectively. The lower panels present the ratio of data to the fits. For the CMS data points, systematic uncertainties are indicated by rectangles (correlated uncertainties) and shaded areas (uncorrelated uncertainties), while vertical solid lines depict statistical uncertainties. In the lower panels, the vertical lines represent the total uncertainties in the data.
Figure 3: Template fits for the inclusive jet $\Delta\text{j}$ distributions in PbPb data (black circles) at 10–30% centrality, measured by the CMS experiment across different $p_{\text{T}}$ intervals. The lower panels show the ratio of data to the fits. For the CMS data points, systematic uncertainties are indicated by rectangles (correlated uncertainties) and shaded areas (uncorrelated uncertainties), while vertical solid lines depict statistical uncertainties. In the lower panels, the vertical lines represent the total uncertainties in the data.
Figure 4: Template fits for the inclusive jet $\Delta\text{j}$ distributions in PbPb data (black circles) at 0–10% centrality, measured by the CMS experiment across different $p_{\text{T}}$ intervals. The lower panels show the ratio of data to the fits. For the CMS data points, systematic uncertainties are indicated by rectangles (correlated uncertainties) and shaded areas (uncorrelated uncertainties), while vertical solid lines depict statistical uncertainties. In the lower panels, the vertical lines represent the total uncertainties in the data.
@@
2.2 Estimation of energy loss via $p_{\text{T}}$ shift for inclusive, quark, and gluon jets
@@
The jets reconstructed experimentally in heavy-ion collisions have been subjected to medium-induced effects (i.e., quenching), and therefore, the final-state reconstructed energy is expected to be less than that of the parton initiating the jet. Since the jet axis decorrelation observable, $\Delta\text{j}$, strongly depends on the initial parton $p_{\text{T}}$, the observed narrowing of the distribution could be interpreted as a momentum shift in the reconstructed jet $p_{\text{T}}$ due to jet quenching. We investigate this hypothesis by performing template fits with systematically shifted pythia templates. The shift is determined by minimizing the $\chi^{2}$ between the $p_{\text{T}}$-shifted $\Delta\text{j}$ distribution in pythia and the CMS PbPb data across the nominal $p_{\text{T}}$ ranges and centrality bins. This procedure involves incrementally increasing the jet $p_{\text{T}}$ by $+1$ GeV in the pythia templates at each step. For instance, for inclusive jets in the PbPb data with $p_{\text{T}}$ in the range $130120<p_{\text{T}}<$ 130 GeV, the corresponding pythia jet $p_{\text{T}}$ range is shifted to $131121<p_{\text{T}}<$ 131 GeV, $132122<p_{\text{T}}<$ 132 GeV, and so on, with the $\chi^{2}$ value computed at each step. A similar method is applied separately to quark and gluon jets using two-dimensional $\chi^{2}$ scans.
The results of this study demonstrate that the average $p_{\text{T}}$ shift required to optimize the data description with pythia templates is largest for the most central 0–10% collisions. The extracted shift values decrease progressively toward more peripheral collisions. In the 50–80% centrality range, the nominal and shifted $p_{\text{T}}$ ranges become nearly identical. Similar trends are observed for fits with templates allowing for the independent variation of quark- and gluon-initiated jets energies, with gluon jet templates consistently preferring larger shift values. The average $p_{\text{T}}$ shift for inclusive jets (i.e., with identical shifts applied to both the quark and gluon templates), as well as for independently shifted quark- and gluon-initiated jet templates, is presented in Table 1 for the 0–10% central collisions, where the greatest energy loss is expected. As anticipated, gluon-initiated jets exhibit a larger $p_{\text{T}}$ shift, consistent with larger energy loss for gluons traversing the medium. These average $p_{\text{T}}$ shifts serve as a quantitative measure of the energy lost by jets during their interactions with the QGP medium.
Nominal $p_{\text{T}}$ (in GeV) $150120<p_{\text{T}}<$ 150 $190150<p_{\text{T}}<$ 190 $230190<p_{\text{T}}<$ 230 $300230<p_{\text{T}}<$ 300
Nominal $p_{\text{T}}\right>$ (in GeV) $131.98\pm 0.016$ $165.50\pm 0.02$ $206.20\pm 0.024$ $255.50\pm 0.036$
Nominal $p_{\text{T}}\right>$ of quark-like jets (in GeV) $132.11\pm 0.022$ $165.68\pm 0.028$ $206.39\pm 0.031$ $255.81\pm 0.047$
Nominal $p_{\text{T}}\right>$ of gluon-like jets (in GeV) $131.87\pm 0.022$ $165.32\pm 0.029$ $205.97\pm 0.035$ $255.05\pm 0.056$
Shifted $p_{\text{T}}$ (in GeV) $172142<p_{\text{T}}<$ 172 $220180<p_{\text{T}}<$ 220 $262222<p_{\text{T}}<$ 262 $345275<p_{\text{T}}<$ 345
Shifted $p_{\text{T}}\right>$ (in GeV) $154.26\pm 0.017$ $196.03\pm 0.023$ $238.58\pm 0.026$ $301.41\pm 0.042$
Shifted $p_{\text{T}}$ of quark-like jets (in GeV) $153123<p_{\text{T}}<$ 153 $190150<p_{\text{T}}<$ 190 $232192<p_{\text{T}}<$ 232 $300230<p_{\text{T}}<$ 300
Shifted $p_{\text{T}}\right>$ of quark-like jets (in GeV) $135.13\pm 0.023$ $165.68\pm 0.028$ $208.41\pm 0.031$ $255.81\pm 0.047$
Shifted $p_{\text{T}}$ of gluon-like jets (in GeV) $168138<p_{\text{T}}<$ 168 $217177<p_{\text{T}}<$ 217 $259219<p_{\text{T}}<$ 259 $347277<p_{\text{T}}<$ 347
Shifted $p_{\text{T}}\right>$ of gluon-like jets (in GeV) $150.12\pm 0.024$ $192.85\pm 0.033$ $235.45\pm 0.039$ $303.05\pm 0.067$
@@
3 Results and discussion

### v2

@@ I Introduction
Ultra-relativistic heavy-ion collision experiments aim to recreate a deconfined state of quarks and gluons, known as the quark-gluon plasma (QGP), under laboratory conditions to investigate the properties of this unique form of nuclear matter. Jets, collimated streams of particles produced in high-energy collisions, serve as versatile tools for probing the QGP [Apolinario:2022vzg, Cunqueiro:2021wls]. These jets are initiated by energetic quarks and gluons (partons) produced in high-momentum-transfer processes during the initial stages of the collision. The partons subsequently fragment into lower-energy partons, which then hadronize into final-state particles. As the outgoing partons traverse the QGP, they lose energy through interactions with the medium. These interactions result in medium-induced modifications to various experimental observables, collectively referred to as “jet quenching”. Experimental evidence for jet quenching has been observed using multiple probes at both the BNL RHIC [PHENIX:2001hpc, STAR:2002ggv] and the CERN LHC [CMS:2021vui, CMS:2021nhn, ATLAS:2023hso, ALICE:2023waz]. Such probes include the suppression of high transverse momentum ($p_{\text{T}}$) particle yields, modifications to jet fragmentation functions, and changes in jet shapes. Some recent reviews on this subject can be found in Refs. [CMS:2024krd, ALICE:2022wpn, Apolinario:2022vzg, Cunqueiro:2021wls].
@@
In parallel with experimental studies of quenching phenomena, significant theoretical advances have been made over the past two decades in characterizing partonic energy loss in the QGP medium. The mechanisms understood as the primary contributors to the energy loss are medium-induced gluon radiation and collisional energy loss. Despite substantial progress on both theoretical and experimental fronts, the details of parton-medium interactions and the precise mechanisms governing parton energy loss remain unclear. Emerging techniques, such as jet substructure studies and the tagging (or selection) of quark and gluon jets, offer new insights into energy loss mechanisms [Kogler:2018hem, Vertesi:2024tdv]. Due to differences in Casimir color factors ($\text{C}_{\text{F}}=4/3$ for quarks and $\text{C}_{\text{A}}=3$ for gluons), gluon-initiated jets are expected to interact more strongly with the QGP and lose more energy compared to quark-initiated jets. This enhanced energy loss reduces the fraction of gluon jets within the surviving jet population in heavy-ion collisions at a given jet $p_{\text{T}}$ compared to the vacuum case without a medium. Recent measurements suggest that the relative fractions of quark and gluon jets are modified in heavy-ion collisions compared to elementary proton-proton collisions, reflecting the differential energy loss due to the distinct color charges of quarks and gluons [mediumqg, eg:partonloss, Zhang:2023oid, ALICE:2019ykw].
@@
The CMS experiment attempted to study gluon-like jet fractions by applying the template-fit method to the jet-charge observable for an inclusive jet sample from lead-lead (PbPb) collisions at a center-of-mass energy per nucleon pair ($\sqrt{s_{{}_{\text{NN}}}}$) of 5.02 TeV [CMS:2020plq]. This measurement found no significant evidence of a decrease in gluon jets or a corresponding increase in quark jets. The jet-charge observable currently lacks theoretical predictions for the expected in-medium modifications. The ATLAS experiment measured the jet nuclear modification factor ($R_{\text{AA}}$) and estimated the fractional energy loss ($S_{\text{loss}}$) for photon-tagged jets, and compared these observables with those for inclusive jets [ATLAS:2023iad]. Since photon-tagged jets are dominantly quark-initiated, this comparison provides sensitivity to the color-charge dependence of parton energy loss [ATLAS:2023iad]. It was reported that the $R_{\text{AA}}$ of photon-tagged jets is significantly higher, and their estimated $S_{\text{loss}}$ significantly lower, than those of inclusive jets, providing evidence that quark jets lose less energy than gluon jets in the medium. However, a detailed description of color-charge dependent energy loss in the QGP medium, along with the relative modifications of quark and gluon jet fractions, remains elusive and warrants further experimental and theoretical investigation through additional observables.
@@
Recently, the CMS experiment measured a new jet substructure observable, the jet axis decorrelation ($\Delta\text{j}$), for a similar data sample (inclusive jets from 5.02 TeV PbPb collisions) [CMS:2024koh]. This observable is predicted to be sensitive to the parton-medium interactions and provide constraints for energy loss models [angelJetaxis]. The study found that $\Delta\text{j}$ is narrower than predictions from various unquenched models, such as pythia [pythia82] and herwig [Sirunyan:2020pqv]. This narrowing could be attributed to modifications in the internal structure of jets caused by interactions with the QGP medium, selection biases favoring less-quenched jets, and/or results from the color-charge dependence of partonic energy loss. In the latter scenario, the larger average energy loss shifts gluon jets further down in energy compared to quark jets, resulting in a different partonic composition of quenched and unquenched jet samples.
@@
In this paper, we attempt to interpret the recent CMS jet axis decorrelation measurements [CMS:2024koh] by comparing them to pythia[pythia82] simulations with varying relative quark/gluon fractions and emulated color-charge dependent jet energy loss. A template-fit approach is employed to estimate the lower limit on gluon jet fractions within the inclusive jet sample. The quark- and gluon-initiated jet templates are derived from inclusive jet samples in 5.02 TeV pp collisions simulated with the pythia8 event generator. Simultaneously, the average jet energy loss for inclusive, quark, and gluon jets is estimated by minimizing the chi-square $(\chi^{2})$ between the $\Delta\text{j}$ distributions measured by CMS and those generated by pythia8 with an applied $p_{\text{T}}$ shift imitating energy loss. Two scenarios are considered. First, in a simplistic approximation, all jets are assumed to lose the same fractional energy, with each jet’s $p_{\text{T}}$ shifted down by the same fraction of transverse momenta. Additionally, we explore the color-charge dependent energy loss scenario, for which quark and gluon jets are allowed independent shifts to account for their differing energy losses.
@@
The paper is organized as follows. Section II outlines the analysis framework used to measure jet axis decorrelation. Section II.1 discusses the template fits, while Section II.2 focuses on estimating the $p_{\text{T}}$ shifts for inclusive, quark, and gluon jets to interpret the CMS $\Delta\text{j}$ distributions. Section III presents the results, followed by a detailed discussion. Finally, Section IV summarizes the main conclusions of this work.
II Analysis method and framework
@@
The jet axis decorrelation, $\Delta\text{j}$, is the angular difference between the axes determined by the anti-$k_{\text{T}}$ energy-weighted (E-scheme) recombination and the winner-take-all (WTA) recombination schemes [angelJetaxis, CMS:2024koh]. It is defined as
$\Delta\mathrm{j}=\sqrt{\bigl(\eta^{\mathrm{E\text{-}scheme}}-\eta^{\mathrm{WTA}}\bigr)^{2}+\bigl(\phi^{\mathrm{E\text{-}scheme}}-\phi^{\mathrm{WTA}}\bigr)^{2}}$ (1)
@@
Where $\eta^{\text{E-scheme}}$ ($\eta^{\text{WTA}}$) and $\phi^{\text{E-scheme}}$ ($\phi^{\text{WTA}}$) denote the pseudorapidity and azimuthal angle of the E-scheme (WTA) axis, respectively, which characterize the jet’s direction. In the E-scheme jet recombination [Cacciari:2011ma], the jet axis is determined by the four-vector sum of its constituent particles, representing a momentum-averaged direction of energy flow that includes contributions from soft particles. This method clusters jets by iteratively combining particle pairs into pseudo-jets, with the direction of each new pseudo-jet defined by the vector sum of the four-momenta of its constituents. The final four-vector defines the E-scheme jet axis in $(\eta,\phi)$ coordinates. Conversely, the WTA scheme emphasizes the hardest prong at each clustering step, making it more sensitive to collinear radiation. To determine the WTA axis, the jet is reclustered using the WTA algorithm [wta1, wta2] with the same constituents, where the direction of each new pseudo-jet aligns with the highest-$p_{\text{T}}$ particle or pseudo-jet. This typically results in the WTA axis aligning with the direction of the hardest constituent. The measured $\Delta\text{j}$ distributions are typically normalized to the total jet number in each $p_{\text{T}}$ interval, constituting a shape-only measurement.
@@
The pythia v8.226 [pythia82] Monte Carlo event generator, configured with the CP5 tune [pythiaCP5tune] and NNPDF3.1 parton distribution functions at next-to-next-to-leading order [nnpdf31], is used to simulate particle production in pp collisions. Following the CMS data analysis, final-state jets are reconstructed using the anti-$k_{\text{T}}$ algorithm [Cacciari:2008gp] with a distance parameter of $R=0.4$, as implemented in the FastJet framework [Cacciari:2011ma]. To match the acceptance of the experimental work, simulated jets are selected within the kinematic range of $120<p_{\text{T}}<300$ GeV and $|\eta|<1.6$.
@@
In pythia, jets can be identified (or tagged) based on the type of parton originating the shower, allowing differentiation between quark and gluon jets. According to pythia8 simulations, the inclusive jet sample in the transverse momentum range of this study consists of comparable contributions from quark and gluon jets in the absence of medium effects, with the gluon-jet fraction being approximately 53% at $p_{\text{T}}=120$ GeV and gradually decreasing to about 45% at $p_{\text{T}}=300$ GeV. The predicted color-charge dependence of partonic energy loss in the presence of QGP is expected to result in a more significant migration of gluon-initiated jets towards lower final-state energy values than quark jets. Since gluon jet $\Delta\text{j}$ distributions are significantly broader than those of quark jets, differences in energy loss (and corresponding migration rates) could also contribute to the narrowing of the jet-axis decorrelations observed in PbPb data by the CMS experiment [CMS:2024koh].
Figure 1: Left: $\Delta R_{\text{axis}}$ distributions for quark and gluon jets in pythia8, illustrating flavor dependence. Right: $p_{\text{T}}$ dependence of the distribution for inclusive jets. The bottom panel shows the ratio of the distributions.
Figure 1 shows $\Delta\text{j}$ distributions from pythia simulations. The left panel compares quark- and gluon-initiated jets in the range $230<p_{\text{T}}<300$ GeV, with quark jets exhibiting narrower $\Delta\text{j}$ distributions than gluon jets. The right panel illustrates the $p_{\text{T}}$ dependence for inclusive jet distributions, comparing jets with $120<p_{\text{T}}<150$ GeV and $230<p_{\text{T}}<300$ GeV, indicating narrowing at higher-$p_{\text{T}}$. In both panels, the distributions exhibit stronger shape dependence at small $\Delta\text{j}$, while at larger $\Delta\text{j}$, there is little to no difference between quark and gluon jets or across different $p_{\text{T}}$ intervals.
II.1 Template fits
@@
A template fit was performed on the CMS PbPb data using distributions associated with different partons derived from pythia simulations. The primary goal was to investigate a potential shift in the rates between quark and gluon jets. Since pythia is a vacuum-like (unquenched) simulation, this fitting attempt is designed to capture the full extent of the observed modifications in the CMS PbPb data by reducing the relative abundance of gluon jets in the measured inclusive jet sample. All quark flavors are combined into a single quark template, and a two-component quark-gluon template fit is applied to each measured CMS distribution. Figure 2 illustrates the template fits for different centrality and jet $p_{\text{T}}$ intervals. Centrality is defined as the fraction of the total nucleus-nucleus cross section, with the interval 0–10% corresponding to the sample with the highest average overlap of the colliding nuclei. The fit range is shown in the $\Delta\text{j}$ distributions, and the lower panels display the ratio of data to fits.
@@
Figure 2: Template fits to the inclusive jet $\Delta\text{j}$ distributions (black circles) in PbPb collisions for 50–80% (top left), 30–50% (top right), 10–30% (bottom left), and 0–10% (bottom right) centrality intervals, measured by CMS across different $p_{\text{T}}$ intervals. Quark and gluon jet templates from pythia are shown as red and blue shaded areas, respectively. The lower panels show the data-to-fit ratios. Systematic uncertainties in CMS data are represented by rectangles (correlated) and shaded areas (uncorrelated), while vertical solid lines denote statistical uncertainties. In the lower panels, vertical lines represent total data uncertainties.
@@
II.2 Estimation of energy loss via $p_{\text{T}}$ shift for inclusive, quark, and gluon jets
@@
The jets reconstructed experimentally in heavy-ion collisions have been subjected to medium-induced effects (i.e., quenching), and therefore, the final-state reconstructed energy is expected to be less than that of the parton initiating the jet. Since the jet axis decorrelation observable, $\Delta\text{j}$, strongly depends on the initial parton $p_{\text{T}}$, the observed narrowing of the distribution could be interpreted as a momentum shift in the reconstructed jet $p_{\text{T}}$ due to jet quenching. We investigate this hypothesis by systematically increasing the jet $p_{\text{T}}$ in pythia to find the best description of the data without varying the expected quark and gluon jet abundances (the fraction of gluon-initiated jets naturally decreases with jet $p_{\text{T}}$ in the kinematic range of this study). The optimal shift is determined by the minimum $\chi^{2}$ between the $p_{\text{T}}$-shifted pythia-based inclusive jet $\Delta\text{j}$ distributions and the CMS PbPb $\Delta\text{j}$ data across the nominal $p_{\text{T}}$ ranges and centrality bins. This procedure involves incrementally increasing the jet $p_{\text{T}}$ by +1+1 GeV in the pythia templates at each step. For instance, for inclusive jets in the PbPb data with $p_{\text{T}}$ in the range $120<p_{\text{T}}<130$ GeV, the corresponding pythia jet $p_{\text{T}}$ range is shifted to $121<p_{\text{T}}<131$ GeV, $122<p_{\text{T}}<132$ GeV, and so on, with the $\chi^{2}$ value computed at each step.
@@
Nominal $p_{\text{T}}$ (in GeV) $120<p_{\text{T}}<150$ $150<p_{\text{T}}<190$ $190<p_{\text{T}}<230$ $230<p_{\text{T}}<300$
Nominal $\left<p_{\text{T}}\right$> (in GeV) $131.98\pm 0.016$ $165.50\pm 0.02$ $206.20\pm 0.024$ $255.50\pm 0.036$
Nominal $\left<p_{\text{T}}\right$> of quark-like jets (in GeV) $132.11\pm 0.022$ $165.68\pm 0.028$ $206.39\pm 0.031$ $255.81\pm 0.047$
Nominal $\left<p_{\text{T}}\right$> of gluon-like jets (in GeV) $131.87\pm 0.022$ $165.32\pm 0.029$ $205.97\pm 0.035$ $255.05\pm 0.056$
Shifted $p_{\text{T}}$ (in GeV) $142<p_{\text{T}}<172$ $180<p_{\text{T}}<220$ $222<p_{\text{T}}<262$ $275<p_{\text{T}}<345$
Shifted $\left<p_{\text{T}}\right$> (in GeV) $154.26\pm 0.017$ $196.03\pm 0.023$ $238.58\pm 0.026$ $301.41\pm 0.042$
Shifted $p_{\text{T}}$ of quark-like jets (in GeV) $123<p_{\text{T}}<153$ $150<p_{\text{T}}<190$ $192<p_{\text{T}}<232$ $230<p_{\text{T}}<300$
Shifted $\left<p_{\text{T}}\right$> of quark-like jets (in GeV) $135.13\pm 0.023$ $165.68\pm 0.028$ $208.41\pm 0.031$ $255.81\pm 0.047$
Shifted $p_{\text{T}}$ of gluon-like jets (in GeV) $138<p_{\text{T}}<168$ $177<p_{\text{T}}<217$ $219<p_{\text{T}}<259$ $277<p_{\text{T}}<347$
Shifted $\left<p_{\text{T}}\right$> of gluon-like jets (in GeV) $150.12\pm 0.024$ $192.85\pm 0.033$ $235.45\pm 0.039$ $303.05\pm 0.067$
The results of this study demonstrate that applying a $p_{\text{T}}$ shift to the momenta of inclusive jets produces a narrowing of the $\Delta\text{j}$ distributions. While the inclusive $p_{\text{T}}$-shift model exhibits greater tension with the data at low $\Delta\text{j}$ compared with the other methods, it provides a consistently better description in the range $\Delta\text{j}\approx 0.03$$0.05$, where the other methods show an apparent tension in the two higher jet $p_{\text{T}}$ bins (see Fig. 3). Nevertheless, because the data uncertainties increase with $\Delta\text{j}$, the inclusive $p_{\text{T}}$-shift model shows substantially worse overall agreement with the data in terms of the $\chi^{2}$ per degree of freedom than that achieved with the two-component fit in central collisions. The average $p_{\text{T}}$ shift values, representing the amount of energy lost by the jets, are found to increase with centrality, reaching the largest values in the most central 0–10% collisions. The extracted shift values decrease progressively toward more peripheral collisions. In the 50–80% centrality range, the nominal and shifted $p_{\text{T}}$ ranges become nearly identical. Similar trends are observed when allowing for the independent variation of quark- and gluon-initiated jet energies, with gluon jets consistently preferring larger $p_{\text{T}}$-shift values. The average $p_{\text{T}}$ shift for inclusive jets (i.e., with identical shifts applied to both the quark and gluon jets), as well as for independently shifted quark- and gluon-initiated jets, is presented in Table 1 for the 0–10% central collisions, where the greatest energy loss is expected. As anticipated, gluon-initiated jets exhibit a larger $p_{\text{T}}$ shift, consistent with larger energy loss for gluons traversing the medium. These average $p_{\text{T}}$ shifts serve as a quantitative measure of the energy lost by jets during their interactions with the QGP medium.
III Results and discussion